what is the substitution solution to x+y=3 and y=x+3
Answer by ikleyn(52784)

.
You just have y expressed via  x  from the 2nd equation,  y = x + 3.
Substitute it into the first equation.  You will get

x + (x+3) = 3.

Hence,  2x = 0, x = 0.

Then  y = x+3 = 0+3 = 3.

Answer.  x = 0, y = 3.